# Simple Fermionic backflow states via a systematically improvable tensor decomposition

### v1

@@ I Introduction
In this work we return to tensor factorizations for a novel wave function parameterization, yet within the framework of NQS. There has been much research to indicate that the prodigious flexibility of complex NQS architectures is not being fully exploited for many correlated problems, due to the challenges in their optimization and initialization within the VMC framework [12, 13]. Many simple parameterizations have performed as accurately as more complex forms, and a premium is placed on the compactness of the ansatz for ease of practical optimization alongside the overall flexibility. The simpler form for these models can also potentially provide tools for easier interpretability and improvability of the many-body physics, and open avenues to alternative optimization strategies. An example of this is the Gaussian process state (GPS), which was originally motivated via Bayesian regression as a systematically improvable kernel model with a single model parameter controlling the expressibility. It was shown to achieve similar quality results to NQS, while often being more compact and open to novel insights [14, 15, 9]. Further development exposed a duality of the GPS wave function model to an exponential of a CANDECOMP/PARAFAC (CP) tensor-rank decomposition of the wave function amplitudes in second quantization [8]. This represented an interesting simplification of the model, and suggests further developments to bring together the domains of NQS and tensor decompositions for systematically improvable descriptions of correlated states.
@@
This is the topic of this work in application to general Fermionic systems, which have proven a particular challenge for NQS methods to date. Arguably the most success for ab initio electronic structure with NQS has come in a first quantized real-space formulation, where permutation-invariant multielectron orbitals are parameterized as a neural network, before their product is explicitly antisymmetrized as a determinant (or short linear combination thereof) [16, 17]. However, in this work we consider a fixed basis and second quantization. While this introduces a (necessarily incomplete) basis set approximation, it also allows for more flexibility in the choice of model (permutational invariance and antisymmetry are automatically enforced). This also allows for problems to be defined in a finite and discrete space for the stochastic sampling where additional approximations can be devised and chemical insights from atomic orbital correlators are easily accessible. Furthermore, this formulation allows for a straightforward treatment of core electrons and direct comparison to established quantum chemical methods, as well as natural application to multi-resolution and quantum embedding methodologies [18, 19, 20]. The basis set approximation, in common with traditional quantum chemical methods, is also much studied with a number of approaches available which can substantially ameliorate it [21, 22, 23, 24].
@@
Here, we consider a particularly simple CP tensor rank decomposition for these configuration-dependent backflow orbitals, which allows for a straightforward yet systematically improvable form for the overall state [46, 47]. We show in Sec. II how this introduces explicit many-body correlated physics whilst retaining antisymmetry of the state. Two previous approaches to combine tensor decompositions with backflow parameterizations have been considered in the literature; a backflow-inspired extension of the MPS ansatz [48] to build non-local entanglement beyond the native MPS tensor ordering constraints for spin systems, as well as a fixed tensor factorization of Fermionic backflow form [42]. The latter study is more relevant to this work, but the particular choice of tensor factorization did not admit a parameter by which the results could be systematically improved, instead relying on a Lanczos step to improve results. In contrast, the CP decomposition of the backflow parameterization introduced in this work overcomes these issues, providing a simple and improvable form which is invariant to orbital ordering.
We also develop a practical approach for ab initio systems to truncate the length scale of the backflow correlations, providing a further compression of the model with little loss of accuracy. We apply this variational ansatz to find the ground state of (doped) Fermi-Hubbard models in Sec. III.1, significantly outperforming a comparable neural network backflow parameterization, as well as considering a challenging ab initio molecular electronic structure problem in Sec. III.2. In common with other studies, we find increasing the sampling of the VMC optimization important to improve results, indicating that despite the simple form of the state it is still challenging to optimize to the expressibility limit of the ansatz [4, 45]. Nevertheless, we still achieve a comparable level of accuracy of more complex parameterizations. In Sec. III.3, we consider a $6\times 6$ 2D lattice of hydrogen atoms as a step towards extended systems, with the CP backflow state comparing favourably to state-of-the-art DMRG calculations and significantly beyond the scope of exact approaches. Finally, in Sec. IV we discuss the computational scaling of the method and approaches to reduce this as an outlook towards larger systems and widespread application.
@@ II Backflow determinants via CP tensor-rank decomposition
The key-idea of backflow ansätze is to extend the Slater determinant by generalizing the single-particle orbitals to functions with non-linear parametric dependencies on all electron coordinates. Historically this meant transforming the electron coordinates $\mathbf{r}_{\alpha}$ with a new set of coordinates $\mathbf{r}^{bf}_{\alpha}=\mathbf{r}_{\alpha}+\sum_{\beta\neq\alpha}\eta(|\mathbf{r}_{\beta}-\mathbf{r}_{\alpha}|)(\mathbf{r}_{\beta}-\mathbf{r}_{\alpha})$, where the function $\eta(\mathbf{r})$ describes the effective displacement of the $\alpha$ electron due to the instantaneous position of the other electrons [49]. This configurational-dependence on all other electron positions can also be directly encoded into the variational parameters of a linear expansion of single-particle orbitals, as first introduced by Ref. [31] for lattice models, yielding a new set of backflow orbitals $\phi^{bf}_{i}\left(\mathbf{r}_{\alpha};\{\mathbf{r}_{/\alpha}\}\right)$. In an effort to improve the systematic description of these configuration-dependent backflow orbitals, recent work has proposed to model $\phi^{bf}_{i}\left(\mathbf{r}_{\alpha};\{\mathbf{r}_{/\alpha}\}\right)$ using neural networks [41, 16, 44], ensuring that the electron labels in $\{\mathbf{r}_{/\alpha}\}$ are invariant under permutation to retain overall antisymmetry.
@@
$\Psi^{\mathrm{CPD}}=\sum_{\mathbf{n}}\Psi^{\mathrm{CPD}}(\mathbf{n})|\mathbf{n}\rangle,$ (7)
@@
Finally, we note that, although the functional form of the configuration-dependent orbitals of Eq. 6 is linear in $M$, it does not reduce to an uncorrelated determinant in the limit of $M=1$. Correlated physics such as that captured via Gutzwiller or Jastrow correlators are included even in this limit, since the dependence between the instantaneous occupation of sites $\mu$ and $\nu$ can be independently addressed in a product form. Indeed, full non-trivial $N$-body correlations are included even at $M=1$, as the exponentially large sum of products of these orbitals formed from the determinant in Eq. 8 builds in an exponential sum of these $N$-fold products of variational parameters for each orbital. This results in a highly-expressive state even for very low $M$, which is systematically improvable to exactness as $M$ is increased.
II.1 Initialization
@@
A practical bottleneck in working with parameterized quantum states with many variational parameters can often be their reliable stochastic optimization. This can be highly sensitive to the initialization of the state, and the simple functional form of the CPD backflow orbitals in Eq. 6 allows for a straightforward and effective initialization of the variational parameters. Specifically, the tensor $\epsilon_{\mu i;n_{\nu}\nu m}$ can be initialized to ensure that the CPD wave function exactly spans a given single determinant such as that found from a prior mean-field solution. For most practical cases, this provides a good starting point for the VMC optimization.
@@
II.2 Backflow truncation via exchange cutoff
@@
Figure 1: Electron density of the orbital subspace chosen via the exchange truncation in the range and rank of explicit backflow correlations. We consider the $K=5$ subspace found for a central atom in a $6\times 6$ lattice of hydrogen atoms in a STO-6G basis at both compressed and extended geometries.
To illustrate the action of this exchange cutoff heuristic, in Figure 1 we consider the electron density of the $K=5$ most entangled orbitals about a specific atom for a $6\times 6$ square grid of ab initio hydrogen atoms in a Boys localized basis [55] at two different interatomic distances, $d$. As expected for a local representation, the exchange cutoff reduces to selection of the nearest-neighbor atomic-localized orbitals around the chosen hydrogen atom in both geometries considered, providing a black-box metric to select the backflow subspace of correlations for each orbital via exploitation of locality of these correlations. We note again that the product structure of the CPD ansatz will build longer-ranged and higher-rank correlations outside the chosen subset implicitly, albeit no longer explicitly for each orbital independently.
@@ III Results
In all results below we initialize the CPD backflow state from the restricted Hartree–Fock solution as outlined in Eq. 9, with a noise scale value $\sigma=0.01$. We optimize parameters using the Stochastic Reconfiguration (SR) method [56], and when the number of parameters is larger than that of the samples, we take advantage of the recently introduced kernel formulation from Ref.  [57] to improve the computational cost of the optimization, as outlined further in Section IV. On the Fermi-Hubbard model, we found that a SR optimizer with RMSProp momentum, as introduced in Ref. [58], outperforms standard SR, and we therefore use this optimizer for the results presented in Section III.1. The final energies presented in the results are computed as averages over 50 independent energy evaluations with the final optimized parameters and a large sample size ($2^{16}$ for the Fermi-Hubbard model and the water molecule, and $2^{14}$ for the $6\times 6$ hydrogen lattice). Error bars are computed as the standard deviation of these independent energy evaluations.
@@ III.1 Fermi-Hubbard model
Figure 2: Percentage relative energy error for the ground state of the $4\times 4$ square Fermi-Hubbard model at $U/t=8$ compared to exact diagonalization results [65] at a hole-doped filling of $n=0.875$ (upper panel) and half-filling (lower panel). CPD backflow results ($\Psi^{CPD}$) are shows as a function of the configurational sample size in the optimization of the parameters, for two different model complexities of $M=1$ and $M=2$. Neural network backflow results are taken from Ref. [41].
@@
We also show the variational improvability as the support dimension $M$ of the CPD decomposition is increased from $M=1$ to $M=2$, with a systematic lowering of all energies found, leading to a maximum of $65,536$ parameters. Nevertheless, we unfortunately find that it is still generally more advantageous to increase the number of configurational samples in the Markov chain than to formally increase the flexibility of the state by increasing $M$. This is due to noise in the estimates of the expectation values required for the optimization of the CPD parameters, which amongst other things affects the inversion of the sampled quantum geometric tensor. This indicates that we cannot be confident of a complete optimization to the global minimum of this state, despite the simple parameterization of the CPD form, with the optimization still limited more by noise in the samples than flexibility in the model, as found in many other studies of comparable states. However, we do find a reliable a systematic improvement in the optimized state as this number of samples is increased, and a high level of accuracy overall for this correlated state.

### v2

@@ I Introduction
In this work we return to tensor factorizations for a novel wave function parameterization, inspired by the developments in the class of NQS descriptions of Fermionic quantum matter. There has been much research to indicate that the prodigious flexibility of complex NQS architectures is not being fully exploited for many correlated problems, due to the challenges in their optimization and initialization within the VMC framework [12, 13]. Many simple parameterizations have performed as accurately as more complex forms, and a premium is placed on the compactness of the ansatz for ease of practical optimization alongside the overall flexibility. The simpler form for these models can also potentially provide tools for easier interpretability and improvability of the many-body physics, and open avenues to alternative optimization strategies. An example of this is the Gaussian process state (GPS), which was originally motivated via Bayesian regression as a systematically improvable kernel model with a single model parameter controlling the expressibility. It was shown to achieve similar quality results to NQS, while often being more compact and open to novel insights [14, 15, 9]. Further development exposed a duality of the GPS wave function model to an exponential of a CANDECOMP/PARAFAC (CP) tensor-rank decomposition of the wave function amplitudes in second quantization [8]. This represented an interesting simplification of the model, and suggests further developments in the use of tensor decompositions for systematically improvable descriptions of correlated states. The potential for synergies between the two domains of variational state parameterizations in NQS and tensor decompositions has in fact already motivated the development of NQS architectures that incorporate MPS parameterizations [16, 17].
@@
This is the topic of this work where we consider a simple and systematically improvable variational quantum state based on tensor factorization, with application to general Fermionic systems which have proven a particular challenge for NQS methods to date. Arguably the most success for ab initio electronic structure with NQS has come in a first quantized real-space formulation, where the single-electron orbitals of a Slater determinant (or short linear combination thereof) have been replaced by multi-electron orbitals parametrized as permutation-invariant neural networks [18, 19, 20]. In this work we consider a fixed basis and second quantization. While this introduces a (necessarily incomplete) basis set approximation, it also allows for more flexibility in the choice of model (permutational invariance and antisymmetry are automatically enforced). This also allows for problems to be defined in a finite and discrete space for the stochastic sampling where additional approximations can be devised and chemical insights from atomic orbital correlators are easily accessible. Furthermore, this formulation allows for a straightforward treatment of core electrons and direct comparison to established quantum chemical methods, as well as natural application to multi-resolution and quantum embedding methodologies [21, 22, 23]. The basis set approximation, in common with traditional quantum chemical methods, is also much studied with a number of approaches available which can substantially ameliorate it [24, 25, 26, 27].
@@
Here, we consider a particularly simple CP tensor rank decomposition for these configuration-dependent backflow orbitals, which allows for a straightforward yet systematically improvable form for the overall state [48, 49]. We show in Sec. II how this introduces explicit many-body correlated physics whilst retaining antisymmetry of the state. Two previous approaches to combine tensor decompositions with backflow parameterizations have been considered in the literature; a backflow-inspired extension of the MPS ansatz [50] to build non-local entanglement beyond the native MPS tensor ordering constraints for spin systems, as well as a fixed tensor representation of two-body Fermionic backflow form [45]. The latter study did not factorize these backflow correlations and was constrained to a two-body form for these correlations. In contrast, the CP decomposition of the backflow parameterization introduced in this work overcomes these issues, providing a simple and improvable form for arbitrary rank correlations which is invariant to orbital ordering.
We also develop a practical approach for ab initio systems to truncate the length scale of the backflow correlations, providing a further compression of the model with little loss of accuracy. We apply this variational ansatz to find the ground state of (doped) Fermi-Hubbard models in Sec. III.1, significantly outperforming a comparable neural network backflow parameterization, as well as considering a challenging ab initio molecular electronic structure problem in Sec. III.2. In common with other studies, we find increasing the sampling of the VMC optimization important to improve results, indicating that despite the simple form of the state it is still challenging to optimize to the expressibility limit of the ansatz [4, 47]. Nevertheless, we still achieve a comparable level of accuracy of more complex parameterizations. In Sec. III.3, we consider a $6\times 6$ 2D lattice of hydrogen atoms as a step towards extended systems, with the CPD backflow state comparing favourably to state-of-the-art DMRG calculations and significantly beyond the scope of exact approaches. Finally, in Sec. IV we discuss the computational scaling of the method and approaches to reduce this as an outlook towards larger systems and widespread application.
@@ II Backflow determinants via CP tensor-rank decomposition
The key-idea of backflow ansätze is to extend the Slater determinant by generalizing the single-particle orbitals to functions with non-linear parametric dependencies on all electron coordinates. Historically this meant transforming the electron coordinates $\mathbf{r}_{\alpha}$ with a new set of coordinates $\mathbf{r}^{bf}_{\alpha}=\mathbf{r}_{\alpha}+\sum_{\beta\neq\alpha}\eta(|\mathbf{r}_{\beta}-\mathbf{r}_{\alpha}|)(\mathbf{r}_{\beta}-\mathbf{r}_{\alpha})$, where the function $\eta(\mathbf{r})$ describes the effective displacement of the $\alpha$ electron due to the instantaneous position of the other electrons [51]. This configurational-dependence on all other electron positions can also be directly encoded into the variational parameters of a linear expansion of single-particle orbitals, as first introduced by Ref. [34] for lattice models, yielding a new set of backflow orbitals $\phi^{bf}_{i}\left(\mathbf{r}_{\alpha};\{\mathbf{r}_{/\alpha}\}\right)$. In an effort to improve the systematic description of these configuration-dependent backflow orbitals, recent work has proposed to model $\phi^{bf}_{i}\left(\mathbf{r}_{\alpha};\{\mathbf{r}_{/\alpha}\}\right)$ using neural networks [44, 18, 20], ensuring that these functions are invariant under permutation of the electron labels in $\{\mathbf{r}_{/\alpha}\}$ to retain overall antisymmetry of the state.
@@
|$\ket{\Psi^{\mathrm{CPD}}}=\sum_{\mathbf{n}}\Psi^{\mathrm{CPD}}(\mathbf{n})|\mathbf{n}\rangle,$ (7)
@@
Finally, we note that, although the functional form of the configuration-dependent orbitals of Eq. 6 is linear in $M$, it does not reduce to an uncorrelated determinant in the limit of $M=1$. Correlated physics such as that captured via Gutzwiller or Jastrow correlators are included even in this limit, since the dependence between the instantaneous occupation of sites $\mu$ and $\nu$ can be independently addressed in a product form. Indeed, full non-trivial $N$-body correlations are included even at $M=1$, as the exponentially large sum of products of these orbitals formed from the determinant in Eq. 8 builds in an exponential sum of these $N$-fold products of variational parameters for each orbital. This results in a highly-expressive state even for very low $M$, which is systematically improvable to exactness as $M$ is increased. We consider the expressibility of these states from an analytic perspective further below.
II.1 Universality of the CPD backflow ansatz and Slater-Jastrow form
The universality of the proposed CPD backflow ansatz directly stems from the universality of the CP decomposition. This means that in the large-$M$ limit, the CP decomposition employed to model the orbitals $\varphi_{\mu i;\mathbf{n}}^{\mathrm{CPD}}$ according to Eq. (6), can be chosen such that they are allowed to vary independently for each many-electron configuration $\mathbf{n}$. This limit also implies a mapping between basis states (Slater determinants) and wavefunction amplitudes after anti-symmetrization which can represent any antisymmetric state without approximation error, according to the definition of the CPD backflow ansatz,
$\Psi^{\mathrm{CPD}}(\mathbf{n})=\mathcal{A}[\varphi^{\mathrm{CPD}}_{\mu_{1}i;\mathbf{n}}\varphi^{\mathrm{CPD}}_{\mu_{2}j;\mathbf{n}}\dots\varphi^{\mathrm{CPD}}_{\mu_{N}k;\mathbf{n}}].$ (9)
This however formally requires $M$ to scale with the number of many-body configurations in the space (as expected for any exact parameterization), as we expand on below.
To show this and make contact with other forms of parameterized states, we consider the subset of CPD states in which the backflow (many-electron) orbitals $\varphi_{\mu i;\mathbf{n}}^{\mathrm{CPD}}$ can factorize into a term which is independent of the specific configuration (an $\mathbf{n}$-independent ‘static’ molecular orbital), and a term which is independent of the site index, but yet can depend on a CP decomposition of the specific many-electron configuration. This can be written as
$\varphi_{\mu i;\mathbf{n}}^{\mathrm{factored-CPD}}=\varphi_{\mu i}\times\left(\sum_{m=1}^{M}\prod_{\nu=1}^{L}\epsilon_{i;n_{\nu}\nu m}\right)$ (10)
With this construction, a product of CP decompositions can be factored out of the determinant, bringing the backflow ansatz into the form of a Slater-Jastrow wavefunction,
$\Psi(\mathbf{n})=$ $\left(\sum_{m=1}^{M}\prod_{\nu=1}^{L}\epsilon_{i;n_{\nu}\nu m}\right)\left(\sum_{m=1}^{M}\prod_{\nu=1}^{L}\epsilon_{j;n_{\nu}\nu m}\right)\cdots$
$\left(\sum_{m=1}^{M}\prod_{\nu=1}^{L}\epsilon_{k;n_{\nu}\nu m}\right)\times\mathcal{A}[\varphi_{\mu_{1}i}\varphi_{\mu_{2}j}\dots\varphi_{\mu_{N}k}].$ (11)
In this, a ‘configuration-universal’ Slater determinant (defined by the $N$ orbitals labelled $i,j,\dots$) is multiplied by a product of $N$ CP decompositions, each of which depends on the local occupations of each site ($n_{\nu}$). This product of CP decompositions takes the place of the Jastrow factor. Assuming that no two rows of the configuration-independent ‘static’ Slater determinant orbital matrix, $\varphi_{\mu i}$, are linearly dependent such that the determinant always evaluates to a non-zero value [36], the universal approximator property of the CP decompositions in the prefactor allows this ansatz to define a one-to-one mapping from each many-electron configuration to an arbitrary wavefunction amplitude.
This approach to factoring out a CPD decomposition from the Slater determinant shows that we formally require a support dimension $M$ which scales as the size of the Hilbert space and is therefore of little practical use. Nonetheless, the representation according to Eq. (II.1) still provides insights into the ability of the ansatz to represent electronic quantum states of interest as a compact model with small support dimension, $M$. From the consideration of the restricted version of the CPD state in this Slater-Jastrow form as shown in Eq. II.1, we find that $M=1$ is sufficient to represent any single Slater determinant, as well as a site-dependent penalty function depending on its local occupation. This encapsulates physically-relevant electronic correlation beyond the mean-field picture. As a specific example, we can consider a parameterization of a Gutzwiller factor of
$\epsilon_{i;n_{\nu}\nu m}=\begin{cases}e^{g_{\nu}}\quad\text{if }i=1\text{ and}n_{\nu}\equiv\uparrow\downarrow\\ 1\quad\text{otherwise}\\ \end{cases},$ (12)
where $n_{\nu}\equiv\uparrow\downarrow$ indicates a double occupancy of the $\nu^{\textrm{th}}$ site [53, 54]. This modulates the Slater determinant with a factor depending on the double occupancy of the sites in each configuration, $\Psi(\mathbf{n})\sim e^{\sum_{\nu}g_{\nu}n_{\nu,\uparrow}n_{\nu,\downarrow}}\times\mathcal{A}[\varphi_{\mu_{1}i}\varphi_{\mu_{2}j}\dots\varphi_{\mu_{N}k}]$, with parameters $g_{\nu}$. General forms for the $\epsilon_{i;n_{\nu}\nu m}$ parameters even at $M=1$ will however also admit factorized non-local dependence on the site occupations beyond Gutzwiller form.
In general, this simple factorization of the CPD state only represents a small subset of the parametrizations possible, which has a significantly larger variational flexibility even for $M=1$. This is because the factorization into a site-dependent term and Slater determinant does not need to be imposed, allowing a non-trivial coupling between these ‘orbital’ and ‘site’ effects at the level of the ansatz. This enlarges the span of states accessible within this decomposition at small $M$, and can therefore outperform ‘Slater-Jastrow’ type factorizations where the Jastrow is taken to have a flexible form, such as those previously considered within the GPS family of states [15].
II.2 Initialization
@@
A practical bottleneck in working with parameterized quantum states with many variational parameters can often be their reliable stochastic optimization. This can be highly sensitive to the initialization of the state, since random initialization of the parameters does not always guarantee a good overlap with the ground state, which can slow down or even prevent the optimization from converging to the true ground state. The simple functional form of the CPD backflow orbitals in Eq. 6 allows for a straightforward and effective initialization of the variational parameters, without the requirement for pre-training [18, 47]. Specifically, the tensor $\epsilon_{\mu i;n_{\nu}\nu m}$ can be initialized to ensure that the CPD wave function exactly spans a given single determinant such as that found from a prior mean-field solution. For most practical cases, this provides a good starting point for the VMC optimization.
@@
II.3 Backflow truncation via exchange cutoff
@@
Figure 1: Electron density of orbital subspaces selected via exchange truncation. The $K=5$ orbital subspace (red) is shown for a central atom in a $6\times 6$ lattice of hydrogen atoms in a STO-6G basis at both (a) compressed and (b) extended geometries.
To illustrate the action of this exchange cutoff heuristic, in Figure 1 we consider the electron density of the $K=5$ most entangled orbitals about a specific atom for a $6\times 6$ square grid of ab initio hydrogen atoms in a Boys localized basis [59] at two different interatomic distances, $d$. This truncation is used later for numerical results in Sec. III.3 to assess the accuracy of the truncation scheme. A choice of $K=5$ respects the local symmetries of each atom, as it enables each atom to be explicitly correlated via the backflow transformations with its four nearest neighbour atoms. As hoped, we find that the exchange cutoff protocol described automatically performs this selection of the nearest-neighbor atomic-localized orbitals around the chosen hydrogen atom in both geometries considered, providing a black-box metric to select the backflow subspace of correlations for each orbital via exploitation of locality of these correlations. We note again that the product structure of the CPD ansatz will build longer-ranged and higher-rank correlations outside the chosen subset implicitly, albeit no longer explicitly for each orbital independently.
@@ III Results
In all results below we initialize the CPD backflow state from the restricted Hartree–Fock solution as outlined in Eq. 13, with a noise scale value $\sigma=0.01$. We optimize parameters using the Stochastic Reconfiguration (SR) method [60], and when the number of parameters is larger than that of the samples, we take advantage of the recently introduced kernel formulation from Ref.  [61] to improve the computational cost of the optimization, as outlined further in Section IV. On the Fermi-Hubbard model and the water molecule, we found that a SR optimizer with RMSProp momentum regularization, as introduced in Ref. [62], outperforms standard SR, and we therefore use this optimizer for the results presented in Section III.1 and III.2. The final energies presented in the results are computed as averages over 50 independent energy evaluations with the final optimized parameters and a large sample size ($2^{16}$ for the Fermi-Hubbard model and the water molecule, and $2^{14}$ for the $6\times 6$ hydrogen lattice). Error bars are computed as the standard deviation of these independent energy evaluations.
@@ III.1 Fermi-Hubbard model
Figure 2: Performance of the CPD and neural network backflow on the Fermi-Hubbard model. Percentage relative energy error for the ground state of the $4\times 4$ square Fermi-Hubbard model at $U/t=8$ compared to exact diagonalization results [69] at (a) a hole-doped filling of $n=0.875$ and (b) half-filling . CPD backflow results ($\Psi^{CPD}$) are shown as a function of the configurational sample size in the optimization of the parameters, for two different model complexities of $M=1$ (blue) and $M=2$ (green). Neural network backflow results (NNB, dashed lines) are taken from Ref. [44].
@@
We also show the variational improvability as the support dimension $M$ of the CPD decomposition is increased from $M=1$ to $M=2$, with a systematic lowering of all energies found, leading to a maximum of $65,536$ parameters. Nevertheless, we unfortunately find that it is still generally more advantageous to increase the number of configurational samples in the Markov chain than to formally increase the flexibility of the state by increasing $M$. This is due to noise in the estimates of the expectation values required for the optimization of the CPD parameters, which amongst other things affects the inversion of the sampled quantum geometric tensor. This indicates that we cannot be confident of a complete optimization to the global minimum of this state, despite the simple parameterization of the CPD form, with the optimization still limited more by noise in the samples than flexibility in the model, as found in many other studies of comparable states. We will consider this behaviour more in the following section, but note that emerging optimization approaches, such as the SPRING algorithm [71], will be able to be transferred to this setting and hold promise to boost the resulting performance of the CPD backflow state. However, despite these current limitations we do find a reliable a systematic improvement in the optimized state as the number of samples is increased, and a high level of accuracy overall for this correlated state.